OpenStudy (anonymous):

find the indefinite integral. (2e^x+3e^(2x))/(e^x)dx i dont know how to integrate this need Help!

OpenStudy (turingtest):

It's all in the simplification:\[\int\limits {(2e^x+3e^{2x})\over(e^x)}dx=\int\limits 2+3e^xdx=2x+3e^x+C\]